If you are riding a bicycle with 27'' tires at a speed of 20 mph, how fast are the wheels rotating?
Answer by Alan3354(69443)

If you are riding a bicycle with 27'' tires at a speed of 20 mph, how fast are the wheels rotating?
-------------
20 mi/hr = 20*22/15 ft/sec = 88/3 ft/sec
Each rev of the tire is 27*pi/12 ft = 2.25*pi ft/rev
(88/3) ft/sec / 2.25 ft/rev = 88/(3*2.25) rev/sec
= 88*60/6.75 rpm
= 782.22 rpm